OpenStudy (anonymous):

Find f(2) if f(x) = (x + 1)2. 5 6 9

OpenStudy (akashdeepdeb):

SUBSTITUTE that is it!! What do YOU think the answer is?

OpenStudy (anonymous):

erm erm erm erm erm erm emr emrme rm6ermemrermermem mr mrm re rmer emr em

OpenStudy (anonymous):

6

OpenStudy (akashdeepdeb):

9 bro! Because you substitute 2 in place of x and then (2+1)^2 = 3^2 = 9!! Understood the concept? Try doing some more on your own!

OpenStudy (akashdeepdeb):

Did you get me?

OpenStudy (anonymous):

i um no